The width of a rectangle is 6 cm less than the length. A second rectangle, with a perimeter of 54 cm, is 3cm wider and 2 cm shorter than the first. Write an equation to help you find the dimensions of both rectangles.
Answer by sofiyac(983)

width of first rectangle = x
length of first rectangle = x+6
width of the second rectangle = x+3
length of second rectangle = x+4
2(x+3)+2(x+4)=54
2x+6+2x+8=54
4x=40
x=10




Substituting x:
width of first rectangle = x = 10 cm
length of first rectangle = x+6 = 16 cm
width of the second rectangle = x+3 = 13 cm
length of second rectangle = x+4 = 14 cm